Crew 181 Daily Summary Report
May 16th, 2017
MDRS Daily Summary Report for SOL 3
Summary Title: Plants, Mice, and EVAs, Oh My!
Mission Status: Quiet day of working on our own projects with a satisfying EVA right before sunset.
Sol Activity Summary: Juan made breakfast of blueberry pancakes for the crew. We all worked on our own tasks for a majority of the day. Charlie worked on checking the water heater, but we still do not have hot water (he will describe more information in his engineering report). The front deck is also sturdy thanks to Charlie’s engineering. We have also been calculating our water usage and feel somewhat confident that we could live another week on our current supply. We seem to be using less than the previous crew, but then again, not many of us have taken showers. A more concrete answer about the next water shipment will be needed in the coming days.
The afternoon was quiet between us all, though we did have a nice discussion about economic sustainability, automation, and how that will affect Mars in the years to come. Avishek planted some seeds in the GreenHab and I recycled some lettuce and chives roots to see if we can perpetuate new growth. We also caught our first Mars vermin, a little mouse, in the Hab kitchen. Janet and Juan worked on drawing and editing their photos and videos.
EVA today consisted of Charlie, Janet, Avishek, and Olly climbing and exploring the Hab Ridge. They had a great time and all arrived back to the Hab safe and sound except for one UV filter on Janet camera (luckily they are inexpensive and she brought a spare.) More information will be in Janet’s journalism report.
Look Ahead Plan: Avishek is making us dinner, then after we will have a bit of crew bonding in the form of a science fiction movie (specific movie – to be determined.) We are requesting an EVA tomorrow using the ATVs for the first time.
Anomalies In Work: None at the present time.
Weather: Sunny, a little cool, wind picked up during the afternoon and died down early during the evening hours. A nice change from last night
Crew Physical Status: All are feeling okay – starting to get into a rhythm of work and sleep.
EVA: Hab Ridge explored by four crew members. Report to follow.
Reports to be Filed: journalist report, engineering report, EVA report, EVA request, and crew patch.
Support Requested: None at the present time.
Prepared by Cassandra
Commander, Crew 181